OpenStudy (anonymous):

Show that when the radius of a circle is double , the angular remains the same and the linear velocity of a point on the circle is doubles.

OpenStudy (amistre64):

Compare circumference: C1 = 2pi R1 C2 = 2pi R2

OpenStudy (amistre64):

R2 = 2 R1 :)

OpenStudy (anonymous):

\[\omega\times r= v\]\[\omega \times 2r =v'\]\[2v=v'\]

OpenStudy (anonymous):

\[2\pi r \frac{\theta}{2\pi}=\text{linear lenght}\] \[ r \theta=\text{linear lenght}\]

OpenStudy (anonymous):

take derivative of both side \[\frac{d\theta}{dt} r=\frac{dl}{dt}\] \[\omega r=v\]

OpenStudy (anonymous):

radius double , than linear velocity is double ?

OpenStudy (anonymous):

if two object A, B have same angular velocity , and A has twice the radius of B, then linear velocity is double for A

OpenStudy (anonymous):

ty